### v3

@@
Compared to MuTAP, a mutation- and LLM-based test generator, CoverUp achieves an overall line+branch coverage of 90% (vs. 77%).
@@
††ccs: Computing methodologies Artificial intelligence††ccs: Software and its engineering Software testing and debugging
@@ 1. Introduction
Compared to MuTAP on one of the benchmark suites it supports, CoverUp again achieves greater or equal coverage on every metric, with an overall line+branch coverage of 90% (vs. 77%).
@@ 4.1. Experimental Setup
Since Codex is no longer available, we use original tests from CodaMosa’s evaluation, using its best performing configuration (0.8-uninterp), available from https://github.com/microsoft/codamosa-dataset;
@@
Since Codex is no longer available, we use original tests from MuTAP’s evaluation, available from https://github.com/ExpertiseModel/MuTAP;
@@
Across the entire benchmark suite, CoverUp achieves 64% (vs. 54% and 51%) line coverage, 49% (vs. 34% and 27%) branch coverage, and 60% (vs. 49% and 45%) line+branch coverage.
On a per-module basis, CoverUp achieves 82% (vs. 63% and 54%) line coverage, 67% (vs. 36% and 19%) branch coverage, and 80% (vs. 56% and 47%) line+branch coverage.
@@
Across the entire benchmark suite, CoverUp achieves 90% (vs. 82% to 16%) line coverage, 89% (vs. 73% to 0%) branch coverage, and 90% (vs. 79% to 10%) line+branch coverage; on a per-module basis, like its baselines, CoverUp achieves 100% on all metrics, except for MuTAP (gtp4o few-shot), which achieves extremely low coverage (17%, 0% and 11%).
@@
Table 2. [RQ1] CoverUp outperforms MuTAP’s coverage on the MT suite.
@@
We observe that 60.3% of successes result from the first prompt, 27.2% from the second, and 12.4% from the third.
@@ C. Tokens
This analysis also suggests potential improvements for CoverUp.
Since even the non-ablated prompt results in some test functions that do not increase coverage, CoverUp should measure the coverage generated by each function individually, rather than for the entire LLM response.
Additionally, CoverUp could omit coverage information when prompting for tests for entirely uncovered functions, thereby reducing the cost of tagging every line in a code segment.
We leave this to future work.
Overall Coverage
Median Per-Module Coverage
Test Generator
Line
Branch
Line + Branch
Line
Branch
Line + Branch
CoverUp (no coverage)
CoverUp (no code context)
CoverUp (no error fixing)
CoverUp (ablated) [RQ2]
Table 5. [RQ5] Effect of ablations on coverage: Removing any of these components from CoverUp lowers its performance.
First Run
Second Run
Third Run
End
Test Generator
InitialCov.
Cov. Incr.Functions
InitialCov.
Cov. Incr.Functions
InitialCov.
Cov. Incr.Functions
Cov.
TotalFunctions
CoverUp
7,366
CoverUp (no coverage)
11,222
Table 6. [RQ5] Importance of coverage-based prompting: As the starting coverage grows with each successive run, the coverage-ablated prompt generates fewer and fewer functions that increase coverage.

### v4

@@
Compared to MuTAP, a mutation- and LLM-based test generator, CoverUp achieves an overall line+branch coverage of 89% (vs. 77%).
@@
††copyright: acmlicensed††doi: 10.1145/3729398††journalyear: 2025††journal: PACMSE††journalvolume: 2††journalnumber: FSE††article: FSE128††publicationmonth: 7††ccs: Computing methodologies Artificial intelligence††ccs: Software and its engineering Software testing and debugging
@@ 1. Introduction
Compared to MuTAP on one of the benchmark suites it supports, CoverUp again achieves greater or equal coverage on every metric, with an overall line+branch coverage of 89% (vs. 77%).
@@ 4.1. Experimental Setup
Since Codex is no longer available, we use original tests from CodaMosa’s evaluation, using its best performing configuration (0.8-uninterp);
@@
Since Codex is no longer available, we use original tests from MuTAP’s evaluation;
@@
Across the entire benchmark suite, CoverUp achieves 64% (vs. 54% and 51%) line coverage, 49% (vs. 34% and 29%) branch coverage, and 60% (vs. 49% and 45%) line+branch coverage.
On a per-module basis, CoverUp achieves 82% (vs. 63% and 54%) line coverage, 67% (vs. 36% and 19%) branch coverage, and 80% (vs. 55% and 47%) line+branch coverage.
@@
Across the entire benchmark suite, CoverUp achieves 89% (vs. 81% to 16%) line coverage, 89% (vs. 73% to 0%) branch coverage, and 89% (vs. 78% to 10%) line+branch coverage; on a per-module basis, like its baselines, CoverUp achieves 100% on all metrics, except for MuTAP (gtp4o few-shot), which achieves extremely low coverage (17%, 0% and 11%).
@@
Table 2. [RQ1] CoverUp outperforms MuTAP’s coverage on the MT suite.
@@
We observe that 60.3% of successes result from the first prompt, 27.2% from the second, and 12.5% from the third.
@@ 6. Conclusion
Altmayer Pizzorno and Berger (2025a)
Juan Altmayer Pizzorno and
Emery D. Berger. 2025a.
plasma-umass/coverup: Automatically
Generating Higher-Coverage Test Suites with AI.
https://doi.org/10.5281/zenodo.15187805
Altmayer Pizzorno and Berger (2025b)
Juan Altmayer Pizzorno and
Emery D. Berger. 2025b.
plasma-umass/pytest-cleanslate: State
pollution handling for CoverUp.
https://doi.org/10.5281/zenodo.15187866